OpenStudy (anonymous):

Simplify the square root of 864.

OpenStudy (anonymous):

\[12\sqrt{6}\] @stamp you have a weak pellet

OpenStudy (anonymous):

864 has three square rootable numbers in it, 9*4*4*6. you can remove the numbers, 9 and both 4's buy square rooting them. you are left with \[3*2*2\sqrt{6} = 12\sqrt{6}\]

OpenStudy (anonymous):

Well, thanks I appreciate that!

OpenStudy (stamp):

8 + 6 + 4 = 18 8 + 6 + 4 is divisible by 9, so 864 is divisible by 9. \[\sqrt{9(96)}=\sqrt{9(3)(32)}=\sqrt{9(3)(16)(2)}=\]\[12\sqrt{6}\]